A certain test preparation course is designed to help students improve their scores on the USMLE exam. A mock exam is given at the beginning and end of the course to determine the effectiveness of the course. The following measurements are the net change in 5 students' scores on the exam after completing the course: 14,11,18,9,19 Using these data, construct a 99% confidence interval for the average net change in a student's score after completing the course. Assume the population is approximately normal.

Step 3 of 4 : Find the critical value that should be used in constructing the confidence interval. Round your answer to three decimal places.
Answer #1

df = n - 1 = 5 - 1 = 4

From T table,

t critical value at 0.01 significance level with 4 df = 4.604